# Hamilton cycles in random digraphs with minimum degree at least one

### v1

@@ 2.3 Phase 2
We can from now on assume that $m\sim\tfrac{1}{2}n\log n$. Let ${\mathcal{C}}_{1}$ be the cycle cover in $D$ corresponding to $M_{1}$. $M_{1}$ being a uniform random perfect matching implies that w.h.p. ${\mathcal{C}}_{1}$ has at most $2\log n$ cycles. In this phase we transform ${\mathcal{C}}_{1}$ into a cycle cover in which each cycle has size at least $n/\log n$. We do this using the edges $E_{2}$. Let $Y_{1}$ be the set of indices $i$ such that $i$ appears at most $\log n/100$ times in the sequence $y_{1},y_{2},\ldots,y_{n\log n/5}$.
@@
Let $W_{i}^{\prime}$ be the number of times $i$ appears in $x_{1},x_{3},\ldots,x_{n\log n/5}$ where x is chosen randomly from $\widehat{\Omega}_{1}$ (see the line following equation (6) above for the definition of $\widehat{\Omega}_{1}$). Let $X_{1}^{\prime}=\left\{W_{i}^{\prime}\leq\log n/100\right\}$. Then, with ${\mathcal{E}}^{\prime}$ as in (6) and where $m_{1}=n\log n/5$ and $L=\log n/100$,
@@
Now, $z\sim\tfrac{1}{2}\log n$ and so
$\sum_{k=1}^{\log n/4}\frac{z^{k}}{k!}\leq 2\left(\frac{e^{1+o(1)}z}{\log n/4}\right)^{\log n/4}\leq n^{0.45}.$ (22)
Furthermore, if $k\geq\log n/4$ then the Chernoff bounds imply that
@@
$\sum_{k=\log n/4}^{10\log n}\frac{z^{k}}{k!}\sum_{r=0}^{L}\binom{k}{r}\left(\frac{2}{5}\right)^{r}\left(\frac{3}{5}\right)^{k-r}\leq\sum_{k\geq\log n/4}\frac{z^{k}}{k!}e^{-k/11}\leq\exp\left\{ze^{-1/11}\right\}\leq n^{0.46}.$ (23)
@@
Explanation: $\binom{n}{\ell}$ counts the number of choices the vertices in our connected set $S$ of size $\ell$. $\binom{2n\log n/15}{\ell-1}$ counts the number of choices for the positions of the $\ell-1$ edges of a tree $T$ contained in $S$ in the ordering of $E_{i}$. There are $\ell^{\ell-2}$ choices for $T$, $2^{\ell}(\ell-1)!$ choices for the ordering of the edges in the sequence y and probability $n^{-2\ell+2}$ that the edges are the ones defining $T$. The large bracketed term bounds the probability that the $\ell$ chosen vertices are in $Y_{2}(L)$.
(c) The probablity that we can find such a cycle can be bounded by
@@
Explanation: we choose the vertices of the cycle $C$ in $\binom{n}{\ell}$ ways and choose a vertex of $C$ in $\ell$ ways and an ordering of the vertices of $C$ in $\ell!$ ways. Then we choose the places of the edges of the cycle in x in $\binom{m}{\ell}$ ways. Then $\left(\frac{\Delta_{1}}{m}\right)^{\ell}$ bounds the probability that these places in the sequence contain the claimed edges. Here $\Delta_{1}=6\log n$ is a bound on the maximum degree in x. If $[\ell]\cap ALTERED=\emptyset$ then we can replace x by y. ∎
@@ 2.4 Elimination of Small Cycles
We now give an informal description of a process which removes a small cycle $C$ from a current PD $\Pi$. We start by choosing an (arbitrary) edge $(v_{0},u_{0})$ of $C$ and then delete it to obtain an NPD $\Gamma_{0}$ with $P_{0}=P(\Gamma_{0})\in{\cal P}(u_{0},v_{0})$, where ${\cal P}(x,y)$ denotes the set of paths from $x$ to $y$ in $D$. The aim of the process is to produce a large set $S$ of NPD’s such that for each $\Gamma\in S$, (i) $P(\Gamma)$ has a least $n_{0}$ edges and (ii) the small cycles of $PD(\Gamma)$ are a subset of the small cycles of $\Pi$. We will show that w.h.p. the endpoints of one of the $P(\Gamma)$’s can be joined by an edge to create a permutation digraph with (at least) one less small cycle.
The basic step in an Out-Phase of this process is to take an NPD $\Gamma$ with $P(\Gamma)$ is a path from $u_{0}$ to $v$ and to examine the edges of the core $C_{2}$ leaving $v$ i.e. the edges going out from the end of the path. Let $w$ be the terminal vertex of such an edge and assume that $\Gamma$ contains an edge $(x,w)$. Then $\Gamma^{\prime}=\Gamma\cup\{(v,w)\}\setminus\{(x,w)\}$ is also an NPD. $\Gamma^{\prime}$ is acceptable if (i) $P(\Gamma^{\prime})$ contains at least $n_{0}$ edges and (ii) any new cycle created (i.e. in $\Gamma^{\prime}$ and not $\Gamma$) also has at least $n_{0}$ edges.
@@
We now increase the the formality of our description. We start Phase 2 with a PD $\Pi_{0}$, say, and a general iteration of Phase 2 starts with a PD $\Pi$ whose small cycles are a subset of those in $\Pi_{0}$. Iterations continue until there are no more small cycles. At the start of an iteration we choose some small cycle $C$ of $\Pi$. There then follows an Out-Phase in which we construct a tree $T_{0}=T_{0}(\Pi,C)$ of NPD’s as follows: the root of $T_{0}$ is $\Gamma_{0}$ which is obtained by deleting an edge $(v_{0},u_{0})$ of $C$.
@@
We now define a set $W$ of used vertices. Initially all vertices in $K_{2}$ are unused i.e. $W=\emptyset$. Whenever we examine an edge $(v,w)$, we add both $v$ and $w$ to $W$. So if $v\not\in W$ then it has at least $\log n/100$ random choices $out(v)$ of neighbours to make in $K_{2}$. Furthermore these choices are almost uniform in that no vertex has a more than $600/n$ chance of being chosen. This is because the in-degrees in $C_{2}$ vary between $\log n/100$ and $\Delta_{1}=6\log n$. We do not allow $|W|$ to exceed $n^{3/4}$.
@@
We consider the construction of our $\nu$ In-Phase trees in two stages. First of all we grow the trees only enforcing condition C (ii) of success and thus allow the formation of small cycles and paths. We try to grow them to depth $t_{0}$. The growth of the $\nu$ trees can naturally be considered to occur simultaneously. Let $L_{i,\ell}$ denote the set of start vertices of the paths associated with the nodes at depth $\ell$ of the $i$’th tree, $i=1,2\ldots,\nu,\ell=0,1,\ldots,t_{0}$. Thus $L_{i,0}=\{u_{0}\}$ for all $i$. We prove inductively that $L_{i,\ell}=L_{1,\ell}$ for all $i,\ell$. In fact if $L_{i,\ell}=L_{1,\ell}$ then the acceptable $E(K_{2})$ edges have the same set of initial vertices and since all of the deleted edges are $E(K_{2})$-edges (enforced by C (ii)) we have $L_{i,\ell+1}=L_{1,\ell+1}$.
@@
Thus with probability 1-$O(n^{-2+o(1)})$ we end up with $\nu$ sets of $\nu$ paths, each of length at least $n/\log^{1/2}n$ where the $i$’th set of paths all terminate in $v_{i}$. The sets $in(v_{i})$ are still unexplored and hence
@@
We begin this phase with a cycle cover of $O(\log^{1/2}n)$ cycles, each of size at least $n/\log^{1/2}n$. We use the edges of $E(K_{3})$ to patch these cycles into a Hamilton cycle. Suppose at some stage of the patching process we have cycles $C_{1},C_{2},\ldots,C_{k},k\geq 2$ where $C_{1}$ is the lexicographically largest cycle and $C_{2}$ is the second lexicographically largest cycle. We look for edges $(i,j)\in C_{1},(k,l)\in C_{2}$ such that $(a_{i},b_{l}),(a_{k},b_{j})\in E(K_{3})$ and then replace $C_{1},C_{2}$ by the cycle $C_{1}+C_{2}+(i,l)+(k,j)-(i,j)-(k,l)$. The probability we cannot find two such edges is at most
@@
This is true regardless of previous patches because all the added edges have one end in $C_{2}$ and the vertices of the second lexicographically largest cycle $C_{2}$ at one stage are disjoint from lexicographically largest cycle at any other stage, the former having been absorbed into $C_{1}$.
@@ 3 Conclusion
As already mentioned, we feel that the above analysis can be extended to get $k$ edge disjoint Hamilton cycles if we condition on minimum degree t least $k$, as long as we have $\sim\frac{1}{2}n\log n$ edges. This is not to say that there might not be unforseen technicalities and so we will not formally claim this.
It is even more challenging to adapt the analysis to the case where we have a linear number of edges and minimum degree $k\geq 2$. This is where the focus of research on this problem should be.
@@
We need to have sharp estimates of the probability that $\sum_{1\leq j\leq n}Z_{j}^{\prime},\sum_{1\leq j\leq n}Z_{j}^{\prime\prime}$ are close to their mean $m$. Let $\sigma$ be as in (5). Section A of [11] proves
@@
(h) Let $X_{j}=Z_{j}^{\prime}(Z_{j}^{\prime}-1)-\mathbb{E}(Z_{j}^{\prime}(Z_{j}^{\prime}-1))=Z_{j}^{\prime}(Z_{j}^{\prime}-1)-mz/n$ for $j\in[n]$. Let $\Delta_{1}=\log^{6}n$.
$\mathbb{P}(X_{j}\geq\Delta_{1})\leq\mathbb{P}(Z_{j}^{\prime}\geq\log^{3}n)=\sum_{j\geq\log^{3}n}\frac{z^{j}}{j!(e^{z}-1)}\leq\left(\frac{e}{\log^{2}n}\right)^{\log^{3}n}.$
Putting $X_{j}^{\prime}=\min\left\{X_{j},\Delta_{1}\right\}$ we see that
@@
$\leq(n^{1/2}\log n)\mathbb{P}\left(\left|\sum_{j=1}^{n}X_{j}^{\prime}\right|\geq n^{2/3}\right)+\mathbb{P}(\exists j:X_{j}\geq\Delta_{1}).$
now $-\log^{2}n\leq X_{j}^{\prime}\leq\Delta_{1}$ and so by Hoeffding’s theorem
$\mathbb{P}(|S_{1}({\bf x})-mz|\geq n^{2/3}/2)\leq(n^{1/2}\log n)\exp\left\{-\frac{n^{4/3}}{2n(\Delta_{1}^{2}+\log^{2}n)}\right\}+n\left(\frac{e}{\log^{2}n}\right)^{\log^{3}n}.$

### v2

@@ 2.3 Phase 2
We can from now on assume that $c_{n}$ does not tend to $\infty$- ∞, using the notation of Lemma 6. There is nothing more to prove if $c_{n}\to-\infty$. We can also assume that $m\leq(1+o(1))n\log n$, for otherwise we can refer to [9]. So now let $m=an\log n$ where $a\in[1/2-o(1),1+o(1)]$, and that $M_{1}$ exists.
Let ${\mathcal{C}}_{1}$ be the cycle cover in $D$ corresponding to $M_{1}$. $M_{1}$ being a uniform random perfect matching implies that w.h.p. ${\mathcal{C}}_{1}$ has at most $2\log n$ cycles. In this phase we transform ${\mathcal{C}}_{1}$ into a cycle cover in which each cycle has size at least $n/\log^{1/2}n$. We do this using the edges $E_{2}$. See Section 2.4 for a description of how this is done.
Let $Y_{1}$ be the set of indices $i$ such that $i$ appears at most $\log n/100$ times in the sequence $y_{1},y_{2},\ldots,y_{n\log n/5}$.
@@
Let $W_{i}^{\prime}$ be the number of times $i$ appears in $x_{1},x_{3},\ldots,x_{n\log n/5}$ where x is chosen randomly from $\widehat{\Omega}_{1}$, where $\widehat{\Omega}_{1}$ is defined in Case 1b, following (16) above. Let $X_{1}^{\prime}=\left\{i:W_{i}^{\prime}\leq\log n/100\right\}$. Then, with ${\mathcal{E}}^{\prime}$ as in (7) and where $m_{1}=n\log n/5$ and $L=\log n/100$,
@@
Now, $z\sim a\log n,a\in[1/2-o(1),1+o(1)]$ (see (6)) and so
$\sum_{k=1}^{a\log n/2}\frac{z^{k}}{k!}\leq 2\left(\frac{2e^{1+o(1)}z}{a\log n}\right)^{a\log n/2}\leq n^{9/10}.$ (18)
Furthermore, if $k\geq a\log n/2$ then the Chernoff bounds imply that
@@
$\sum_{k=a\log n/2}^{10\log n}\frac{z^{k}}{k!}\sum_{r=0}^{L}\binom{k}{r}\left(\frac{2}{5}\right)^{r}\left(\frac{3}{5}\right)^{k-r}\leq\sum_{k\geq a\log n/2}\frac{z^{k}}{k!}e^{-k/11}\leq\exp\left\{ze^{-1/11}\right\}\leq n^{0.46}.$ (19)
@@
Explanation: $\binom{n}{\ell}$ counts the number of choices for the vertices in our connected set $S$ of size $\ell$. $\binom{2n\log n/15}{\ell-1}$ counts the number of choices for the positions of the $\ell-1$ edges of a tree $T$ contained in $S$ in the ordering of $E_{i}$. There are $\ell^{\ell-2}$ choices for $T$, $2^{\ell}(\ell-1)!$ choices for the ordering of the edges in the sequence y and probability $n^{-2\ell+2}$ that the edges are the ones defining $T$. The large bracketed term bounds the probability that the $\ell$ chosen vertices are in $Y_{2}(L)$.
(c) The probability that we can find such a cycle can be bounded by
@@
Explanation: we choose the vertices of the cycle $C$ in $\binom{n}{\ell}$ ways and choose a vertex of $C$ in $\ell$ ways and an ordering of the vertices of $C$ in $\ell!$ ways. Then we choose the places of the edges of the cycle in x in $\binom{m}{\ell}$ ways. Then $\left(\frac{\Delta_{1}}{m}\right)^{\ell}$ bounds the probability that these places in the sequence contain the claimed edges. recall that $\Delta_{1}=6\log n$ is a bound on the maximum degree in x. If $C\cap ALTERED=\emptyset$ then we can replace x by y. ∎
@@ 2.4 Elimination of Small Cycles
We now give an informal description of a process which removes a small cycle $C$ from a current PD $\Pi$. We start by choosing an (arbitrary) edge $(v_{0},u_{0})$ of $C$ and then delete it to obtain an NPD $\Gamma_{0}$ with $P_{0}=P(\Gamma_{0})\in{\cal P}(u_{0},v_{0})$, where ${\cal P}(x,y)$ denotes the set of paths from $x$ to $y$ in $D$. The aim of the process is to produce a large set $S$ of NPD’s such that for each $\Gamma\in S$, (i) $P(\Gamma)$ has at least $n_{0}$ edges and (ii) the small cycles of $PD(\Gamma)$ are a subset of the small cycles of $\Pi$. We will show that w.h.p. the endpoints of one of the $P(\Gamma)$’s can be joined by an edge to create a permutation digraph with (at least) one less small cycle.
The Out-Phase consists of a sequence of basic steps. In a basic step of the Out-Phase we have an NPD $\Gamma$ where $P(\Gamma)$ is a path from $u_{0}$ to another vertex $v$. We examine the edges of the core $C_{2}$ leaving $v$ i.e. the edges going out from the end of the path. Let $w$ be the terminal vertex of such an edge and assume that $\Gamma$ contains an edge $(x,w)$. Then $\Gamma^{\prime}=\Gamma\cup\{(v,w)\}\setminus\{(x,w)\}$ is also an NPD. $\Gamma^{\prime}$ is acceptable if (i) $P(\Gamma^{\prime})$ contains at least $n_{0}$ edges and (ii) any new cycle created (i.e. in $\Gamma^{\prime}$ and not $\Gamma$) also has at least $n_{0}$ edges.
@@
We now increase the formality of our description. We start Phase 2 with a PD $\Pi_{0}$, say, and a general iteration of Phase 2 starts with a PD $\Pi$ whose small cycles are a subset of those in $\Pi_{0}$. Iterations continue until there are no more small cycles. At the start of an iteration we choose some small cycle $C$ of $\Pi$. There then follows an Out-Phase in which we construct a tree $T_{0}=T_{0}(\Pi,C)$ of NPD’s as follows: the root of $T_{0}$ is $\Gamma_{0}$ which is obtained by deleting an edge $(v_{0},u_{0})$ of $C$.
@@
We denote the the $E_{i}$-out-neighbors of $v\in K_{2}$ by $out(v)$. We define a set $W$ of used vertices. Initially all vertices in $K_{2}$ are unused i.e. $W=\emptyset$. Whenever we examine an edge $(v,w)$, we add both $v$ and $w$ to $W$. So if $v\not\in W$ then it has at least $\log n/100$ random choices $out(v)$ of neighbours to make in $K_{2}$. Furthermore these choices are almost uniform in that no vertex has a more than $600/n$ chance of being chosen. This is because the in-degrees in $C_{2}$ vary between $\log n/100$ and $\Delta_{1}=6\log n$. We do not allow $|W|$ to exceed $n^{3/4}$.
@@
We consider the construction of our $\nu$ In-Phase trees in two stages. First of all we grow the trees only enforcing condition C (ii) of success and thus allow the formation of small cycles and paths. We try to grow them to depth $t_{0}$. The growth of the $\nu$ trees can naturally be considered to occur simultaneously. Let $L_{i,\ell}$ denote the set of start vertices of the paths associated with the nodes at depth $\ell$ of the $i$’th tree, $i=1,2\ldots,\nu,\ell=0,1,\ldots,t_{0}$. Thus $L_{i,0}=\{u_{0}\}$ for all $i$. We prove inductively that $L_{i,\ell}=L_{1,\ell}$ for all $i,\ell$. In fact if $L_{i,\ell}=L_{1,\ell}$ then the acceptable $E(K_{2})$ edges have the same set of initial vertices and since all of the deleted edges are $E(K_{2})$-edges (enforced by C (ii)) we have $L_{i,\ell+1}=L_{1,\ell+1}$. The fact that $L_{i,\ell}=L_{1,\ell}$ gives us some control over the set of new vertices exposed in the In-Phase.
@@
Thus with probability 1-$O(n^{-2+o(1)})$ we end up with $\nu$ sets of $\nu$ paths, each of length at least $n/\log^{1/2}n$ where the $i$’th set of paths all terminate in $v_{i}$. The sets $in(v_{i})$ have not yet been exposed by the process and hence
@@
We begin this phase with a cycle cover of $O(\log^{1/2}n)$ cycles, each of size at least $n/\log^{1/2}n$. We use the edges of $E(C_{3})$ to patch these cycles into a Hamilton cycle. Suppose at some stage of the patching process we have cycles $L_{1},L_{2},\ldots,L_{k},k\geq 2$ where $L_{1}$ is a largest cycle and $L_{2}$ is a second largest cycle. We look for edges $(i,j)\in L_{1},(k,l)\in L_{2}$ such that $(a_{i},b_{l}),(a_{k},b_{j})\in E(C_{3})$ and then replace $L_{1},L_{2}$ by the cycle $L_{1}+L_{2}+(i,l)+(k,j)-(i,j)-(k,l)$. Given that the degrees of vertices in $C_{3}$ are at least $\log n/100$, the probability we cannot find two such edges is at most
@@
This is true regardless of previous patches because all the added edges have one end in $L_{2}$ and the vertices of the second lexicographically largest cycle $L_{2}$ at one stage are disjoint from lexicographically largest cycle at any other stage, the former having been absorbed into $L_{1}$.
@@ 3 Conclusion
As already mentioned, we feel that the above analysis can be extended to get $k$ edge disjoint Hamilton cycles if we condition on minimum in- and out-degree at least $k$, as long as we have $\sim\frac{1}{2}n\log n$ edges. This is not to say that there might not be unforeseen technicalities and so we will not formally claim this.
It is even more challenging to adapt the analysis to the case where we have a linear number of edges and minimum degree $k\geq 2$. This is where the focus of research on this problem should now be.
@@
We need to have sharp estimates of the probability that $\sum_{1\leq j\leq n}Z_{j}^{\prime},\sum_{1\leq j\leq n}Z_{j}^{\prime\prime}$ are close to their mean $m$. Let $\sigma$ be as in (5). Section A of [11] proves the following proposition:
Proposition 1.
@@
(h) Let $X_{j}=Z_{j}^{\prime}(Z_{j}^{\prime}-1)-\mathbb{E}(Z_{j}^{\prime}(Z_{j}^{\prime}-1))=Z_{j}^{\prime}(Z_{j}^{\prime}-1)-mz/n$ for $j\in[n]$. Let $A=\log^{6}n$.
$\mathbb{P}(X_{j}\geq A)\leq\mathbb{P}(Z_{j}^{\prime}\geq\log^{3}n)=\sum_{j\geq\log^{3}n}\frac{z^{j}}{j!(e^{z}-1)}\leq\left(\frac{e}{\log^{2}n}\right)^{\log^{3}n}.$
Putting $X_{j}^{\prime}=\min\left\{X_{j},A\right\}$ we see that
@@
$\leq(n^{1/2}\log n)\mathbb{P}\left(\left|\sum_{j=1}^{n}X_{j}^{\prime}\right|\geq n^{2/3}\right)+\mathbb{P}(\exists j:X_{j}\geq A).$
Now $-\log^{2}n\leq X_{j}^{\prime}\leq A$ and so by Hoeffding’s theorem
$\mathbb{P}(|S_{1}({\bf x})-mz|\geq n^{2/3}/2)\leq(n^{1/2}\log n)\exp\left\{-\frac{n^{4/3}}{2n(A^{2}+\log^{2}n)}\right\}+n\left(\frac{e}{\log^{2}n}\right)^{\log^{3}n}.$